# Hydrodynamics of Relativistic Superheated Bubbles

### v1

@@ 5 Efficiency factor
6 Large jump in the number of degrees of freedom
@@ 7 Discussion
We have found two qualitative differences between superheated and supercooled flows. The first one is that some superheated detonations and hybrids can develop a necessarily metastable region behind the bubble wall, which at sufficiently long times will decay by nucleating bubbles of the $L$-phase. The second difference is that the pressure inside an expanding superheated bubble can be higher or lower than the pressure outside it. Even in the latter case, the bubble still expands by accreting energy from the outside. This raises an interesting dynamical question about how this steady state is reached from the slightly over-critical bubble that is initially nucleated. The reason is that, for the critical bubble, the inside pressure must be slightly higher than the outside pressure, since it must exactly compensate for the effect of the latter plus the surface tension. It would be interesting to see the time evolution of the inside pressure along the first instants after nucleation in a microscopic model, following the analysis in the supercooled case Bea:2021zsu . An additional insight provided by such a model would be a prediction of the bubble wall velocity from the microscopic dynamics. We will report on these aspects elsewhere.
@@ Appendix A Non-relativistic, superheated bubbles with an incompressible phase
$-{\bf\nabla}\left(\frac{p}{\rho_{L}}\right)=\frac{\partial\bf{v}}{\partial t}+\left({\bf v}\cdot{\bf\nabla}\right){\bf v}-\nu\nabla^{2}{\bf v}$ (32)

### v2

@@ V Efficiency factor
VI Large jump in the number of degrees of freedom
@@ VII Discussion
We have found two qualitative differences between superheated and supercooled flows. The first one is that some superheated detonations and hybrids can develop a necessarily metastable region behind the bubble wall, which at sufficiently long times will decay by nucleating bubbles of the $L$-phase. The second difference is that the pressure inside an expanding superheated bubble (at $\xi=0$) can be higher or lower than the pressure asymptotically far away from the bubble (at $\xi=1$). Even in the latter case, the bubble still expands by accreting energy from the outside. This raises an interesting dynamical question about how this steady state is reached from the slightly over-critical bubble that is initially nucleated. The reason is that, for the critical bubble, the inside pressure must be slightly higher than the outside pressure, since it must exactly compensate for the effect of the latter plus the surface tension. It would be interesting to see the time evolution of the inside pressure along the first instants after nucleation in a microscopic model, following the analysis in the supercooled case Bea:2021zsu . An additional insight provided by such a model would be a prediction of the bubble wall velocity from the microscopic dynamics. We will report on these aspects elsewhere.
@@ Appendix A Non-relativistic, superheated bubbles with an incompressible phase
${\bf\nabla}\left(\frac{p}{\rho_{L}}\right)=\frac{\partial\bf{v}}{\partial t}+\left({\bf v}\cdot{\bf\nabla}\right){\bf v}-\nu\nabla^{2}{\bf v}$ (40)